How do you put (2,4) and (4,-2) in function notation form?
Answer by Theo(13342)

f(2) = 4
f(4) = -2

for example:

suppose you have a function that takes the value of x and squares it.

the function would be f(x) = x^2

now if you evaluate that function for the value of x = 2, you will get:

f(2) = 2^2 which becomes f(2) = 4

now if you have a function that takes the the square value of x and then multiplies it by -1, then that function would be:

f(x) = -sqrt(x)

now if you evaluate that function for when x = 4, you would get:

f(4) = -sqrt(4) which becomes f(4) = -2